Af Reference Card

For af version 2.4 development Copyright \begin{matrix}c \\ ◯ \end{matrix}1996 - 2002 Malc Arnold

Starting Af

To start af in interactive mode, just type “af”. To read a folder instead of your incoming mailbox use “af -f folder”. To start af to send a single message use “af address”.

Leaving Af

suspend af  C-z
exit af  C-x C-c
save all buffers and exit af  M-z

Aborting Commands

Typing “C-g” will always abort a partiallytyped command. If you are in Minibuffer or Typeout mode, “C-g” with no partially-typed command will cancel the operation, returning you to Mail mode.

Getting Help

To get into af’s help system, type “C-h”. To see what options are available, type “C-h” again. For a verbose listing of the options, type “C-h” a third time.

view the af user manual  C-h i
show what matches a string  C-h a
show the help for a function  C-h f
show the help for a key  C-h k
show what command a key runs C-h c

Moving Around the Buffer

move to previous message  C-p or p
move to next message  C-n or n
move to start of buffer  C-< or <
move to end of buffer  C-> or >
move to centre of window  M-.
move to a given line  M-g
scroll forward a screen  C-v
scroll backward a screen  M-v
redraw screen with point centred C-l

Handling Messages

open and read a message  C-o or RET
open the “raw” form of a message C-x C-o
display a message via a pager  C-M-o or

M-RET

open body parts into a new buffer C-x j

open body parts in the other window

C-x 4 j

show info about a message  C-x =
save a message to a folder  M-+
print a message  M-p
pipe a message into a command M-|
edit a message  C-x C-e

Composing MIME Messages

attach a file  a
compose a body part  c
edit a body part  e
edit a body part’s description  d
edit message headers  h
finish composing message  q

Sending Mail

send a mail message  M-s or M-m
send a file as a message  M-a
reply to the current message  M-r
group reply to the message  C-M-r
forward the current message  M-f
bounce the current messsage  M-b
attach the current message  C-M-f
digestify the region  C-x C-d
digestify the tagset  C-t C-d
set a mail alias  C-x C-a

Killing and Yanking

kill the current message  C-k
kill the messages in the region  C-w
copy the region as if killed  M-w
kill tagged messages  C-t C-k
copy tagged messages as if killed C-t w
yank the last kill into the buffer C-y
cycle yanks through kill ring  M-y

Typeout

scroll forward or exit at end  SPC
scroll forward one screen  C-v
scroll backward one screen  M-v
scroll forward one line  C-n
scroll backward one line  C-p
move to start of buffer  C-< or <
move to end of buffer  C-> or >
redraw screen with point centred C-l
search typeout buffer forwards  C-s
search typeout buffer backwards C-r

The Minibuffer

move forwards one character  C-f
move backwards one character  C-b
move forwards one word  M-f
move backwards one word  M-b
move to beginning of line  C-a
move to end of line  C-e
delete the previous character  DEL
delete the next character  C-d
delete to end of line  C-k
delete the previous word  M-DEL
delete the next word  M-d
transpose chars around point  C-t
set the mark  C-SPC
delete the region  C-w
copy region as if killed  M-w
Make the next word lower case  M-l
Make the next word upper case M-u
Make the next word capitalised M-c
Move to the previous history line C-p
Move to the next history line  C-n
Move to the first history line  M-<
Move to the last history line  M->
Search history backwards  C-r
Search history forwards  C-s
quoted insert a character  C-q
complete the current line  TAB
complete the current word  SPC
show possible completions  ?
finish minibuffer input  RET

Af Reference Card

For af version 2.4 development

Mark & Region

set the mark  C-SPC
swap point and mark  C-x C-x
save messages in region  C-x +
print messages in region  C-x p
pipe messages in region  C-x |

Tags

tag the current message  C-t t
untag the current message  C-t u
show tags on current message  C-t ?
save tagged messages  C-t +
print tagged messages  C-t p
pipe tagged messages  C-t |

Searching

search forwards for regex  C-s
search backwards for regex  C-r
search forwards for tag  C-t C-s
search backwards for tag  C-t C-r
tag messages matching regex  C-t s

Sorting

To sort the messages in the buffer use “M-x sort-buffer”. Typing “?” at the prompt will list the sort orders af supports. You can also use “M-x sort-region” or “M-x sort-tagset” to sort the region or tagged messages.

Narrowing

narrow buffer to the region  C-x n
narrow buffer to tagged messages C-t n
widen the buffer  C-x w

Folders

find a folder  C-x C-f
find a folder read-only  C-x C-r
find an alternate folder  C-x C-v
find folder in other window  C-x 4 b
resynchronise the buffer  C-x r
save a buffer to disk  C-x C-s
save some buffers to disk  C-x s
write buffer to a named file  C-x C-w

Buffers

select another buffer  C-x b
select buffer in other window  C-x 4 b
list the buffers  C-x C-b
kill a buffer  C-x k

Windows

split the window in two vertically C-x 2
move to another window  C-x o
scroll the other window  C-M-v
delete the selected window  C-x 0
delete all other windows  C-x 1
enlarge the selected window  C-x ∧